Question 5 Tries remaining:2 C Points out of 6.25(Do not include a $ sign in your response. Round to the nearest two decimal places if necessary.) PFlag question Suppose a firm has an average total cost function given by: ATC 9610/q +1 +10q Calculate the price where the firm breaks even (le profit so Answer: Check
Answer #1

At the break-even price, ATC is minimized and Price = ATC.

ATC is minimized when dATC/dq = 0.

dATC/dq = - (9610 / q2) + 10 = 0

9610 / q2 = 10

q2 = 961

q = 31

Break-even price = ATCminimum = (9610/31) + 1 + (10 x 31) = 310 + 1 + 310 = 621